Solve the system:
4x - 2y + 2z = 38
3y - 5z = -37
4z = 20

Answer by checkley75(3666)

4Z=20, Z=20/4, Z=5 ANSWER.
3Y-5*5=-37, 3Y-25=-37, 3Y=-37+25, 3Y=-12, Y=-12/3, Y=-4 ANSWER.
4X-2*-4+2*5=38, 4X+8+10=38, 4X=38-8-10, 4X=20, X=20/4, X=5 ANSWER.